OpenStudy (anonymous):

How do you find exact value? Cos (- 3pi/4)

OpenStudy (michele_laino):

hint: we can write this: \[\begin{gathered} \cos \left( { - \frac{{3\pi }}{4}} \right) = \cos \left( {\frac{\pi }{4} - \pi } \right) = \hfill \\ \hfill \\ = \cos \left( {\frac{\pi }{4}} \right)\cos \left( \pi \right) + \sin \left( {\frac{\pi }{4}} \right)\sin \left( \pi \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

What happened to the three?

OpenStudy (michele_laino):

we have: \[\frac{\pi }{4} - \pi = - \frac{{3\pi }}{4}\]

OpenStudy (dayakar):

cos (-theta) = cos theta

OpenStudy (dayakar):

cos(-3pi/4)= cos (3pi/4)

OpenStudy (anonymous):

How does that end up as - 1/√ 2?

OpenStudy (dayakar):

cos(3*180/4)= cos(3*45) = cos135

OpenStudy (dayakar):

cos(90+45) = -sin45

OpenStudy (dayakar):

=-1/square root of 2

OpenStudy (anonymous):

Wow. That was way simpler than I made it out to be. Thanks a bunch!